numerical methods question. please explain/solve clearly that is understandable. will give thumbs up if solved in 30 mins correctly

Find the LU decomposition of the matrix. 25 5 4 10 8 16 8 12 22
Answer #1

let 5 25 Ang 4 2. 16 10 8. 12 22 To finding LU - decomposition findo very easy to For that take ALU ; where L is lower a tria

8 K, Ч У А) 2 = 25 - 2 Э ч, 19,9 = y - ua : -1,, ,, 22 - - 16х5 25 Чэр = 6 ДЧ, «Х, Ч, 2 = 12 = 12 ), , , la 29. р 1 12 - 2х5

Ugg 25 L and into U Now, all but Jet 25 5 4 25 5 6 4 360 25 106 25 8 16 22 25 2 52 8 12 25